OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –5/2 (x + 4)

OpenStudy (katrinakaif):

Slope intercept form: y =mx +b y - 7 = -5/2 (x + 4) y - 7 = -5/2x -10 y = -5/2x - 10 (+7) y = -5/2x -3